Two point charges are fixed on the y axis: a negative point charge q1 = -29 μC at y1 = +0.16 m and a positive point charge q2 at y2 = +0.34 m. A third point charge q = +8.3 μC is fixed at the origin. The net electrostatic force exerted on the charge q by the other two charges has a magnitude of 24 N and points in the +y direction. Determine the magnitude of q2.
Answer #1

given that

q1 = -29*10^-6 C

q = 8.3*10^-6 C

y1 = 0.16 m

y2 = 0.34 m

force f = 24 N

electrostatic force between two point charges is f = kq1q2/r^2

here

f net = f1+f2

24 = kqq1/y1^2-kqq2/y2^2

k = 9*10^9

kqq2/y2^2 = 9*10^9*8.3*10^-6*29*10^-6/0.16^2-24

kqq2/y2^2 = 84.62-24 = 60.62 N

q2 = 60.62*0.34^2/8.3*10^-6*9*10^9

q2 = 0.09381*10^-3 C

q2 = 93.81*10^-6 C = 93.81 μC